# Multi-times Monte Carlo Rendering for Inter-reflection Reconstruction

### v2

@@ Abstract
Inverse rendering methods have achieved remarkable performance in reconstructing high-fidelity 3D objects with disentangled geometries, materials, and environmental light. However, they still face huge challenges in reflective surface reconstruction. Although recent methods model the light trace to learn specularity, the ignorance of indirect illumination makes it hard to handle inter-reflections among multiple smooth objects. In this work, we propose Ref-MC2 that introduces the multi-times Monte Carlo sampling which comprehensively computes the environmental illumination and meanwhile considers the reflective light from object surfaces. To address the computation challenge as the times of Monte Carlo sampling grow, we propose a specularity-adaptive sampling strategy, significantly reducing the computational complexity. Besides the computational resource, higher geometry accuracy is also required because geometric errors accumulate multiple times. Therefore, we further introduce a reflection-aware surface model to initialize the geometry and refine it during inverse rendering. We construct a challenging dataset containing scenes with multiple objects and inter-reflections. Experiments show that our method outperforms other inverse rendering methods on various object groups. We also show downstream applications, e.g., relighting and material editing, to illustrate the disentanglement ability of our method. Our project page: https://zhutengjie.github.io/Ref-MC2/
@@ 1 Introduction
Previous methods [27, 9, 43, 54] have explored how to disentangle geometry, diffuse, roughness, metalness, and environmental light from multi-view images, but they still face challenges in shading and reflective objects. Specifically, Nvdiffrec [27] proposes a differentiable pipeline that enables gradient-based optimization on both meshes and volumetric textures. These methods ignore the shadow when modeling the illumination, resulting in failed disentanglement for diffuse and shading appearances. To model more realistic shadings, Nvdiffrecmc [9] further incorporates ray tracing and Monte Carlo sampling [26] into inverse rendering, significantly improving the decomposition of shape, materials, and lighting. However, it ignores the indirect illuminations during path tracing and treats rays attacking object surfaces as sheltered illumination. As a consequence, methods like Nvdiffrecmc increase the ambiguity of geometry reconstruction and undermine material learning. Unfortunately, the scene with multiple inter-reflections is common in the real world. The failure in these scenes hinders the wider applications of these methods. Recent methods [17, 38] take inter-reflections into account and can reconstruct the reflective objects well. However, the implicit representations of these methods for materials and renders sacrifice their scalability to downstream tasks. Recently, Nefii [43] further incorporated the implicit neural radiance to estimate ray tracing, alleviating the ambiguity between materials and indirect illuminations. However, this leads to a huge computational consumption for path tracing.
@@
To address these challenges, we correspondingly propose two strategies. 1) For efficiency, we flatten the multi-times sampling into sequential single-time sampling. In a Lambert model [30], the diffuse light is independent of the direction of reflection and thus can be pre-computed as a map to query at any time. When we trace the reflected light from an object, instead of recalculating the diffuse component, we can take the value directly from the diffuse map. Besides, we also limit the specular component to a small lobe along with the reflective direction. 2) For geometry, we refer to the SDF-based methods [39, 17, 6] and replace the common positional encoding with the Sphere Gaussian encoding to get an accurate initial geometry for reflective objects. We use this geometry to initialize Flxicubes  [33] that optimizes surface meshes based on the gradient, and further fine-tunes the Flexicubes in a differentiable pipeline. To evaluate our framework, we construct a dataset containing difficult scenes which contain strong inter-reflections between multiple surfaces. Extensive experiments demonstrate that our framework can successfully decompose indirect illumination and materials. In summary, our contributions are:
@@
  • We reduce the computational consumption when tracing indirect illumination by pre-computing the diffuse map based on the Lambert model.
@@ 2.2 Neural Inverse Rendering
Although neural implicit surfaces have achieved impressive performance in geometry reconstructing and novel views synthesis, they do not obtain the fundamental materials of PBR which limits their flexibility in downstream tasks. Neural inverse rendering methods [17, 44, 8, 2, 51, 54, 13, 36, 21] introduce the physical rendering equation to estimate the disentangled diffuse and specular component from RGB images. They approximate the rendering equation based on neural networks or basis functions, e.g., Spherical Gaussians [40, 45, 47, 51] and Spherical Harmonics [5, 1, 35, 49]. Nvdiffrec [27] introduces the full inverse rendering that estimates shape, materials, and environmental light into gradient-based optimization. However, it does not consider the shadows, leading to the entanglement of materials. Recent works [6, 39] extend SDF-based architectures with an additional appearance branch to model the reflections on the object. RefNeuS [6] introduces a reparametrization method to distinguish the reflective appearance from the diffuse appearance by a direction-relative process, but it fails to faithfully reconstruct non-reflective objects. UniSDF [39] proposes to use a weight-MLP to balance the reflective and non-reflective branches for different objects. However, these methods still face challenges in scenes with complex inter-reflections. Nvdiffrecmc [9] extends Nvdiffrec with Monte Carlo sampling to trace the light path but still ignores indirect illumination between objects. Further methods [17, 18, 43] consider indirect illumination in their design. ENVIDR [17] employs a neural renderer to learn the physical light interaction, without explicitly formulating the rendering equation. NeRO [18] applies the split-sum approximation to approximate the shading effects of both direct and indirect lights. Nefii [43] introduces ray tracing to the radiance field to model indirect illuminations. However, these methods do not fully disentangle the materials from RGB images, but decompose the appearance into reflective and diffuse color. It is hard to apply to the current industrial pipeline directly. In contrast, our Ref-MC2 is a full inverse rendering method that also considers indirect illumination.
@@ 3.1 Preliminaries
where $L_{i}$($\mathbf{p}$, $\bm{\omega}_{i}$) is the incoming radiance into $\mathbf{p}$ from the direction $\bm{\omega}_{i}$, $\mathbf{n}$ is the normal of the point $\mathbf{p}$, $\Omega$ is the hemisphere of directions above $\mathbf{p}$, $f(\mathbf{p},\bm{\omega}_{i},\bm{\omega}_{o})$ is the BSDF evaluated for $\bm{\omega}_{i}$ and the current incoming direction $\bm{\omega}_{i}$. A physics-based BSDF function, GGX [23] is proposed to decompose the function into several physical terms. The GGX is described as:
@@
Figure 1: We perform Monte Carlo sampling at the viewpoint. When the sampling ray from the point is not blocked, it is the direct illumination from environmental lighting. When the sampling ray hits an object, we divide this indirect illumination from the object into diffuse light and specular light. For diffuse light, we pre-compute the coarse diffuse color under one-time sampling and store it as a diffuse map to query during the multi-times sampling. For specular light, we only need to partially trace the rays in a small specular lobe along the reflective direction. The gradients are backward along the tracing path, and are passed to optimize $\bm{k}_{d},\bm{k}_{orm}$, normals, and environment maps.
@@ 3.2 Multi-times Monte Carlo Sampling
In this equation, $r\left(\mathbf{p},\bm{\omega}_{o}\right)$ represents the location of a surface point on the object surface hit by a ray cast from $\mathbf{p}$ in direction $\bm{\omega}_{i}$ for the first time. The corresponding Monte Carlo integration that considers the indirect illumination can then be expressed as:
$L_{o}\left(\mathbf{p},\bm{\omega}_{o}\right)\approx\frac{1}{N}\sum_{i=1}^{N}\frac{L_{i}\left(r\left(\mathbf{p},\bm{\omega}_{o}\right),-\bm{\omega}_{i}\right)f\left(\mathbf{p},\bm{\omega}_{i},\bm{\omega}_{o}\right)\left(\bm{\omega}_{i}\cdot\mathbf{n}\right)}{p\left(\bm{\omega}_{i}\right)}.$ (6)
@@
In GGX, the diffuse term $f_{d}$ is:
@@
where $c_{\text{diff}}$ is the diffuse albedo of the material. In this equation, $f$ is related to the direction of the incoming radiance, which can yield more realistic results. However, it brings significant computations and limits the flexibility during Monte Carlo integration. Because the term $f_{d}$ is related to the normal $n$ and the direction of rays ${\omega}_{o}$, we cannot pre-compute the diffuse map before deep ray tracing. The following term is to introduce the roughness for diffuse light to avoid too dark edges in extremely low grazing angles. However, in our task, the common sampling hardly faces such a problem, while this term limits the Flexible computation in inverse rendering. To reduce the burden, according to  [24] we can use the Lambert diffuse to approximate it as:
$f_{d}=\frac{c_{\text{diff}}}{\pi}.$ (9)
In this approximation, $f_{d}$ is independent on the incoming direction $\omega_{i}$ and the outgoing direction $\omega_{o}$. Based on this, the diffuse part of outgoing light can be simplified as:
@@
It no longer relates to the direction of the sampling ray, so we can pre-compute it via an MLP $M$. Compared to the diffuse part, the specular part is strongly related to the ray direction. It is a disadvantage that we cannot approximate the specular part like the diffuse part, but it is also an advantage that we only need to sample minor rays in a small specular lobe along the reflective directions. Therefore, we significantly reduce the computations of both the diffuse part and the specular part during the multi-times Monte Carlo Sampling.
@@ 3.3 Geometry Initialization
We first utilize SDF-based architectures [46, 6, 18] to learn an initial geometry. The two branches of the diffuse and reflective networks well disambiguate the appearance with reflections and empower the SDF network to produce a high-quality geometry, as shown in Fig. 2 (a). However, this architecture encounters challenges of indirect illumination due to the expressive capacity of Integrated Positional Encoding (IPE). It performs well in general scenes but shows limitations in representing inter-reflections. Due to the capacity of Sphere Gaussians (SG) to represent radiance directions, we introduce SG encoding instead of IPE to enhance the expressive capacity of reflective MLP, as shown in Fig. 2 (b). Besides, we directly parameterize diffuse appearance as SG coefficients which is suitable for objects with multiple reflective surfaces. The geometry comparison is also shown in Fig. 2.
@@ 4.1 Implementation Details
Method Nvdiffrec [27] Nvdiffrecmc [9] Nefii [43] Ours Ours (w/o Acc.) Ours (w/o Geo.) Ours (3) Ours (1)
@@ 4.3 Multi-times Monte Carlo Sampling
4.4 Initial Geometry
In this section, we further study the necessity of the initial geometry discussed in Sec. 3.3. As shown in Fig. 5, without initial geometry, the rendering results show bumpy surfaces with badly learned materials in the area of hollow shape. Because multi-times sampling amplifies the errors introduced by geometry, the learned materials are even worse than the single-time sampling. When applying a good-quality initialization, the network bypasses the ambiguity brought by the geometry. It helps our method learn well-disentangled materials and finally produces high-quality rendering results. Quantitative comparison in Tab. 1 also demonstrate the necessity of geometry initialization.
@@ 4.5 Relighting and Material Editing
The disentangled environment map and material empower our methods to relight and edit reconstructed objects in the downstream application. Our method can also be seamlessly plugged into the industrial pipeline. As shown in Fig. 6, our method enables the arbitrary combination of reconstructed objects. We can easily edit their metalness, roughness, and albedo color by manipulating the learned materials. Due to the well-disentangled shading, our rendering results are natural and realistic in all five environment maps, even the point light in a dark environment. They also perform well after editing materials thanks to the well-learned material map. For example, after we increase the metalness and reduce the roughness of the teapot, the teapot clearly reflects the neighboring objects on its surface. In another case, when we extremely increase the reflectance and change the base color of the toaster, it can totally reflect the scene and other objects.
@@ 5 Conclusion and Limitations
In conclusion, our Ref-MC2 introduces multi-times Monte Carlo sampling into the inverse rendering pipeline to model the indirect illumination. It improves the performance in scenes with complex inter-reflections. However, increment of sampling times significantly increases computational consumption and makes the pipeline highly geometry-sensitive. To solve the challenge of computational efficiency, based on the Lambert model, we pre-compute the diffuse light that is irrelevant to the viewing direction as a map and directly query it during multi-times sampling. To improve the geometry quality, we adopt SDF-based architecture to get an initial geometry and refine a design with Spherical Gaussian encoding for reflective objects. We further use Flexicubes to take the initial mesh into the differentiable rendering pipeline that learns disentangled materials. Our Ref-MC2 still has several limitations. The major limitation is that the 2-time sampling cannot handle extremely reflective objects, e.g., mirrors, as the specular energy hardly degrades after reflections. Besides, the training time needs to further reduce in future time.

### v3

@@ Abstract
Inverse rendering methods have achieved remarkable performance in reconstructing high-fidelity 3D objects with disentangled geometries, materials, and environmental light. However, they still face huge challenges in reflective surface reconstruction. Although recent methods model the light trace to learn specularity, the ignorance of indirect illumination makes it hard to handle inter-reflections among multiple smooth objects. In this work, we propose Ref-MC2 that introduces the multi-times Monte Carlo sampling which comprehensively computes the environmental illumination and meanwhile considers the reflective light from object surfaces. To address the computation challenge as the times of Monte Carlo sampling grow, we propose a specularity-adaptive sampling strategy, significantly reducing the computational complexity. Besides the computational resource, higher geometry accuracy is also required because geometric errors accumulate multiple times. Therefore, we further introduce a reflection-aware surface model to initialize the geometry and refine it during inverse rendering. We construct a challenging dataset containing scenes with multiple objects and inter-reflections. Experiments show that our method outperforms other inverse rendering methods on various object groups. We also show downstream applications, e.g., relighting and material editing, to illustrate the disentanglement ability of our method. Our project page: https://zhutengjie.github.io/Ref-MC2/.
@@ 1 Introduction
Previous methods [28, 9, 45, 57] have explored how to disentangle geometry, diffuse, roughness, metalness, and environmental light from multi-view images, but they still face challenges in shading and reflective objects. Specifically, Nvdiffrec [28] proposes a differentiable pipeline that enables gradient-based optimization on both meshes and volumetric textures. These methods ignore the shadow when modeling the illumination, resulting in failed disentanglement for diffuse and shading appearances. To model more realistic shadings, Nvdiffrecmc [9] further incorporates ray tracing and Monte Carlo sampling [27] into inverse rendering, significantly improving the decomposition of shape, materials, and lighting. However, it ignores the indirect illuminations during path tracing and treats rays attacking object surfaces as sheltered illumination. As a consequence, methods like Nvdiffrecmc increase the ambiguity of geometry reconstruction and undermine material learning. Unfortunately, the scene with multiple inter-reflections is common in the real world. The failure in these scenes hinders the wider applications of these methods. Recent methods [17, 40, 22] take inter-reflections into account and can reconstruct the reflective objects well. However, the implicit representations of these methods for materials and renders sacrifice their scalability to downstream tasks. Recently, Nefii [45] further incorporated the implicit neural radiance to estimate ray tracing, alleviating the ambiguity between materials and indirect illuminations. However, this leads to a huge computational consumption for path tracing. Neural Microfacet Fields [22] employs two-bounce sampling for indirect light modeling, but its geometry field limits flexibility for downstream tasks.
@@
To address these challenges, we correspondingly propose two strategies. 1) For efficiency, we flatten the multi-times sampling into sequential single-time sampling. In a Lambert model [31], the diffuse light is independent of the direction of reflection and thus can be presented as a map to query at any time. When we trace the indirect light from an object, instead of recalculating the diffuse component, we can take the value directly from the diffuse map. It can be optimized through self-supervision. Therefore, we only need to sample the specular component within a small lobe along around the reflective direction. 2) For geometry, we refer to the SDF-based methods [41, 17, 6] and replace the common positional encoding with the Sphere Gaussian encoding to get an accurate initial geometry for reflective objects. We use this geometry to initialize Flxicubes  [35] that optimizes surface meshes based on the gradient, and further fine-tunes the Flexicubes in a differentiable pipeline. To evaluate our framework, we construct a dataset containing difficult scenes which contain strong inter-reflections between multiple surfaces. Extensive experiments demonstrate that our framework can successfully decompose indirect illumination and materials. In summary, our contributions are:
@@
  • We reduce the computational consumption when tracing indirect illumination by self-supervising the diffuse map based on the Lambert model.
@@ 2.2 Neural Inverse Rendering
Although neural implicit surfaces have achieved impressive performance in geometry reconstructing and novel views synthesis, they do not obtain the fundamental materials of PBR which limits their flexibility in downstream tasks. Neural inverse rendering methods [17, 46, 8, 2, 54, 57, 13, 38, 21] introduce the physical rendering equation to estimate the disentangled diffuse and specular component from RGB images. They approximate the rendering equation based on neural networks or basis functions, e.g., Spherical Gaussians [42, 47, 49, 54] and Spherical Harmonics [5, 1, 37, 52]. Nvdiffrec [28] introduces the full inverse rendering that estimates shape, materials, and environmental light into gradient-based optimization. However, it does not consider the shadows, leading to the entanglement of materials. Recent works [6, 41] extend SDF-based architectures with an additional appearance branch to model the reflections on the object. RefNeuS [6] introduces a reparametrization method to distinguish the reflective appearance from the diffuse appearance by a direction-relative process, but it fails to faithfully reconstruct non-reflective objects. UniSDF [41] proposes to use a weight-MLP to balance the reflective and non-reflective branches for different objects. However, these methods still face challenges in scenes with complex inter-reflections. Nvdiffrecmc [9] extends Nvdiffrec with Monte Carlo sampling to trace the light path but still ignores indirect illumination between objects. Further methods [17, 18, 45] consider indirect illumination in their design. ENVIDR [17] employs a neural renderer to learn the physical light interaction, without explicitly formulating the rendering equation. NeRO [18] applies the split-sum approximation to approximate the shading effects of both direct and indirect lights. Nefii [45] introduces ray tracing to the radiance field to model indirect illuminations. Neural Microfacet Fields [22] employs two-bounce sampling to accurately calculate indirect illumination. However, NMF represents geometry by a density field, which limits the scalability and flexibility for downstream tasks. In addition, these aforementioned methods do not fully disentangle the materials from RGB images, but decompose the appearance into reflective and diffuse color. It is hard to apply to the current industrial pipeline directly.eq: rendering equation In contrast, our Ref-MC2 is a full inverse rendering method that also considers indirect illumination.
@@ 3.1 Preliminaries
where $L_{i}$($\mathbf{p}$, $\bm{\omega}_{i}$) is the incoming radiance into $\mathbf{p}$ from the direction $\bm{\omega}_{i}$, $\mathbf{n}$ is the normal of the point $\mathbf{p}$, $\Omega$ is the hemisphere of directions above $\mathbf{p}$, $f(\mathbf{p},\bm{\omega}_{i},\bm{\omega}_{o})$ is the BSDF evaluated for $\bm{\omega}_{i}$ and the current incoming direction $\bm{\omega}_{i}$. The GGX [24] physics-based BSDF function is proposed to decompose the function into several physical terms. The function is described as:
@@
Figure 1: We perform Monte Carlo sampling at the viewpoint. When the sampling ray from the point is not blocked, it is the direct illumination from environmental lighting. When the sampling ray hits an object, we divide this indirect illumination from the object into diffuse light and specular light. We sample the diffuse light from a diffuse map that is optimized through self-supervision. For specular light, we only need to partially trace the rays in a small specular lobe along the reflective direction. The gradients are backward along the tracing path, and are passed to optimize $\bm{k}_{d},\bm{k}_{orm}$, normals, and environment maps.
@@ 3.2 Multi-times Monte Carlo Sampling
In this equation, $r\left(\mathbf{p},\bm{\omega}_{i}\right)$ represents the location of a surface point on the object surface hit by a ray cast from $\mathbf{p}$ in direction $\bm{\omega}_{i}$ for the first time. The corresponding Monte Carlo integration that considers the indirect illumination can then be expressed as:
$L_{o}\left(\mathbf{p},\bm{\omega}_{o}\right)\approx\frac{1}{N}\sum_{i=1}^{N}\frac{L_{i}\left(r\left(\mathbf{p},\bm{\omega}_{i}\right),-\bm{\omega}_{i}\right)f\left(\mathbf{p},\bm{\omega}_{i},\bm{\omega}_{o}\right)\left(\bm{\omega}_{i}\cdot\mathbf{n}\right)}{p\left(\bm{\omega}_{i}\right)}.$ (6)
@@
In Disney PBR equation [24], the diffuse term $f_{d}$ is:
@@
where $c_{\text{diff}}$ is the diffuse albedo of the material. In this equation, $f$ is related to the direction of the incoming radiance, which can yield more realistic results. However, using this equation to calculate the diffuse light for indirect illumination creates unnecessary extra computation. Because the term $f_{d}$ is related to the normal $n$ and the direction of rays ${\omega}_{o}$, we cannot get the diffuse map before deep ray tracing. The following term is to introduce the roughness for diffuse light to avoid too dark edges in extremely low grazing angles. However, using this equation to compute the diffuse light for indirect illumination introduces unnecessary extra computation. To reduce the burden, we can follow H and approximate it using Lambertian diffuse lighting.:
$f_{d}^{\text{ind}}=\frac{c_{\text{diff}}}{\pi}.$ (9)
Where $f_{d}^{\text{ind}}$ is the diffuse light of the indirect illumination. In this approximation, $f_{d}^{\text{ind}}$ is independent on the incoming direction $\omega_{i}$ and the outgoing direction $\omega_{o}$. Based on this, the diffuse part of indirect light can be simplified as:
@@
where $L_{\text{diff}}^{\text{ind}}$ is the diffuse part of the indirect illumination. It no longer relates to the direction of the sampling ray, so we can present it via an MLP $M$. Compared to the diffuse part, the specular part is strongly related to the ray direction. It is a disadvantage that we cannot approximate the specular part like the diffuse part, but it is also an advantage that we only need to sample minor rays in a small specular lobe along the reflective directions. Therefore, we significantly reduce the computations of both the diffuse part and the specular part during the multi-times Monte Carlo Sampling.
@@ 3.3 Geometry Initialization
We first utilize SDF-based architectures [48, 6, 18] to learn an initial geometry. The two branches of the diffuse and reflective networks well disambiguate the appearance with reflections and empower the SDF network to produce a high-quality geometry, as shown in Fig. 2 (a). However, this architecture encounters challenges of indirect illumination due to the expressive capacity of Integrated Positional Encoding (IPE). It performs well in general scenes but shows limitations in representing inter-reflections. Due to the capacity of Spheical Gaussians (SG) [50, 33] to represent radiance directions, we introduce SG encoding instead of IPE to enhance the expressive capacity of reflective MLP, as shown in Fig. 2 (b). Besides, we directly parameterize diffuse appearance as SG coefficients which is suitable for objects with multiple reflective surfaces. The geometry comparison is shown in Fig. 2.
@@ 4.1 Implementation Details
Method NDR [28] NDRMC [9] Nefii [45] Ours Ours (w/o Acc.) Ours (w/o Geo.) Ours (3) Ours (1)
@@ 4.3 Multi-times Monte Carlo Sampling
4.4 Geometry
In this section, we further explore the necessity of the initial geometry discussed in Sec. 3.3 and present our geometry reconstruction results. As shown in Fig. 5, without initial geometry, the rendering results show bumpy surfaces with badly learned materials in the area of hollow shape. Because multi-times sampling amplifies the errors introduced by geometry, the learned materials are even worse than the single-time sampling. When applying a good-quality initialization, the network bypasses the ambiguity brought by the geometry. It helps our method learn well-disentangled materials and finally produces high-quality rendering results. Quantitative comparison in Tab. 1 also demonstrate the necessity of geometry initialization. In addition, we showcase a real-scene geometry reconstruction result of the NeRO [18] dataset in Fig. 6. We further use Chamfer Distance to quantitatively evaluate our geometric quality and present the results in Tab. 2.
Figure 6: Real scene geometry comparison. We compare our reconstructed geometry with NeRO, Nvdiffrec and Nvdiffrecmc. Our reconstructed geometry demonstrates superior results.
Table 2: Geometry qualitative comparison.We use the Chamfer Distance ($\downarrow$) to evaluate our reconstructed geometry. As shown in the table, we obtain the best results in multiple scenes.
Dataset Nvdiffrec [28] Nvdiffrecmc [9] NeRO [18] Ours
Materials 0.016 0.016 0.0057 0.0030
Coral 0.28 0.25 0.13 0.13
@@ 4.4 Geometry
Figure 8: Relighting results of real scenes. Our method demonstrates more realistic results under several lighting conditions than NeRO.
@@ 4.5 Relighting and Material Editing
The disentangled environment map and material empower our methods to relight and edit reconstructed objects in the downstream application. Our method can also be seamlessly plugged into the industrial pipeline. As shown in Fig. 7, our method enables the arbitrary combination of reconstructed objects. We can easily edit their metalness, roughness, and albedo color by manipulating the learned materials. Due to the well-disentangled shading, our rendering results are natural and realistic in all five environment maps, even the point light in a dark environment. They also perform well after editing materials thanks to the well-learned material map. For example, after we increase the metalness and reduce the roughness of the teapot, the teapot clearly reflects the neighboring objects on its surface. In another case, where we extremely increase the reflectance and change the base color of the toaster, it accomplishes to reflect the scene and other objects. To showcase further applications of our method, we conduct experiments on real-world datasets captured by NeRO [18] and present the comparison results in Fig. 8. It can be seen that Ref-MC2 achieves more realistic results under different lighting conditions.
@@ 5 Conclusion and Limitations
In conclusion, our Ref-MC2 introduces multi-times Monte Carlo sampling into the inverse rendering pipeline to model the indirect illumination. It improves the performance in scenes with complex inter-reflections. However, increment of sampling times significantly increases computational consumption and makes the pipeline highly geometry-sensitive. To solve the challenge of computational efficiency, based on the Lambert model, we simplify the BRDF for indirect lighting, which allows us to reduce the number of ray traces. To improve the geometry quality, we adopt SDF-based architecture to get an initial geometry and refine a design with Spherical Gaussian encoding for reflective objects. We further use Flexicubes to take the initial mesh into the differentiable rendering pipeline that learns disentangled materials. Our Ref-MC2 still has several limitations. The major limitation is that the 2-time sampling cannot handle extremely reflective objects, e.g., mirrors, as the specular energy hardly degrades after reflections. Besides, the training time needs to further reduce in future time.
6 Acknowledgements
This work was supported by NSFC (62201342) and Shanghai Science and Technology Major Project (2021SHZDZX0102). We also thank Student Innovation Center of SJTU for providing GPUs.